• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 02:09:55 on May10,2011

    Entry number 358768

    This entry is a followup to: 358750

    keyword=L-Q1 attempted repair, etc.

    Alex and Bob went into the hall and tried to reset L-Q1.
    We also tried the manual controls in the hall.

    It came up to any desired current but always the wrong polarity.
    Of course, Howard already tried earlier, so I didn't really
    think we could outdo him. But we tried.

    Right now we are checking the R-HRS timing issue. Then we may
    try to run a few hours with L-Q1 off and the consequent poor
    acceptance, for tests.